OpenStudy (anonymous):

Find the volume of a cone with radius 30 in and slant height 50 in.

OpenStudy (anonymous):

r = 30 in slant ht , l= 50in first find height h h ^2 = l^2 - r^2 h^2 = 2500 - 900 h^2 = 1600 h = 40in volume V= pi r^2 h /3 => V= pi x 900 x 40 /3 => V = 1200 pi in ^3

OpenStudy (anonymous):

A. 3769.91 in3 B. 3768 in3 C. 113040 in3 D. 37699.11 in3

OpenStudy (anonymous):

dont know sorry